# Basic representations of genus zero nonabelian Hodge spaces

### v2

@@
Abstract.
@@
1. Introduction
1.1. General motivation: classification of (wild) nonabelian Hodge spaces
@@
The term ‘nonabelian Hodge space’, introduced in [10, Def. 7], refers to a class of manifolds that can be seen as moduli spaces for different types of objects. On one hand, they are moduli spaces of meromorphic connections with possibly irregular singularities on curves (the de Rham side). Via the Riemann-Hilbert-Birkhoff correpondence, they are also moduli spaces of generalised monodromy data, known as wild character varieties [8, 13] (the Betti side), and, via the nonabelian Hodge correspondence [4], as moduli spaces of irregular Higgs bundles (the Dolbeault side). Furthermore, they carry rich (symplectic, hyperkähler) geometric structures, and give rise to (isospectral and isomonodromic) integrable systems, including many classical integrable systems of interest in mathematical physics such as all Painlevé equations (see e.g. [10] for an overview).
@@
A nonabelian Hodge space $\mathcal{M}$ depends on the choice of a tuple $(\Sigma,\mathbf{a},\bm{\Theta},\bm{\mathcal{C}})$, that we call a wild Riemann surface with boundary data. Here the triple $(\Sigma,\mathbf{a},\bm{\Theta})$ is a wild Riemann surface [8, 15], meaning that $\Sigma$ is a compact Riemann surface, $\mathbf{a}=\{a_{1},\dots,a_{m}\}\subset\Sigma$ is a finite set of points on $\Sigma$ corresponding to the singularities of the connections giving rise to points of $\mathcal{M}_{DR}$, the global irregular class $\bm{\Theta}$ encodes the type of singularities that these connections have at each singular point, and $\bm{\mathcal{C}}$ is a collection of conjugacy classes, encoding the formal monodromies of the connections.
@@
Remarkably, it turns out that there exist some isomorphisms between moduli spaces defined by different wild Riemann surfaces with boundary data, corresponding to connections with different ranks, number of singularities and pole orders. While for most known examples such isomorphisms have not been worked out on all the three sides, it is expected that when this happens the moduli spaces are really isomorphic as nonabelian Hodge spaces, i.e. all the structures of the nonabelian Hodge package match in the natural way. In other words, we would like to view those different wild Riemann surfaces with boundary data as different realizations, or representations, of one single abstract nonabelian Hodge space $\mathcal{M}$, which raises the question of their classification. Notice that here we will only be considering representations coming from the case of vector bundles, however wild nonabelian Hodge spaces arising from vector bundles may also be isomorphic to nonabelian Hodge spaces arising from principal $G$-bundles, for $G$ a complex reductive group different from $\operatorname{GL}_{n}(\mathbb{C})$, see e.g. [12].
@@
Varying the wild Riemann surface, one can consider moduli spaces of irregular connections in families, giving rise to isomonodromic deformations [8]. If the connections are written down explicitly in coordinates, this gives systems of nonlinear PDEs which include, among the simplest nontrivial examples, all Painlevé equations [33]. In this picture, the wild Riemann surface $(\Sigma,\mathbf{a},\bm{\Theta})$ giving rise to a given isomonodromy system corresponds to what is called a Lax representation, Lax pair, or linearisation for the system in the literature on integrable systems, albeit in a more abstract sense than is often the case (typically the term Lax pair is used to refer to an explicit parametrisation in coordinates of the coefficients of the connections). The existence of nontrivial isomorphisms between different nonabelian Hodge spaces is related to the existence of different Lax representations for many Painlevé-type equations, and the question of the classification of nonabelian Hodge spaces extends to the problem of finding all possible Lax representations for a given isomonodromy system.
1.2. Diagrams and isomorphisms between moduli spaces of connections
@@
Interestingly, it turns out that in many cases, there is a clean combinatorial way to pass between different representations.
Possibly the simplest instance where such “dualities” between different moduli spaces of connections occur in genus zero is Harnad duality [1, 28, 47, 2], which relates dual connections on trivial vector bundles over the affine line of the form
$d-(A+P(Y-z)^{-1}Q)dz,\qquad\text{and}\qquad d+(Y+Q(A-z)^{-1}P)dz,$
@@
where $P:W\to V$ and $Q:V\to W$ are linear maps, where $V,W$ are finite dimensional complex vector spaces, and $A\in\operatorname{End}(V)$, $Y\in\operatorname{End}(W)$. The connections on both sides have an irregular singularity at infinity, with a pole of order two, as well as regular singularities at finite distance, whose positions are given by the eigenvalues of $A$ and $Y$. Harnad showed that the isomonodromic deformations of these two connections, governed by the JMMS equations [32], are equivalent. The duality induces an isomorphism between the de Rham moduli spaces corresponding to both sides.
@@
Harnad duality can be conveniently described in a graphical way as follows. The linear maps $A,Y,P,Q$ can be viewed as representations of the quiver drawn below, and the duality corresponds to exchanging the roles of the two vertices.
$V$$W$$A$$Y$$P$$Q$
@@
This multiduality can also be conveniently encoded in a graphical way, in terms of a supernova quiver: it consists of a core, which is a complete $k$-partite graph for some integer $k$ (i.e. the core is a disjoint union of $k$ subsets of vertices, and each vertex in one subset is linked by one edge to every vertex belonging to a different subset) to which are then glued legs, each leg being a linear (i.e. type A) quiver.
The main statement is then as follows [7]: such a supernova quiver, together with some extra data, canonically determines $k$+1 moduli spaces of connections of simply-laced type, with in general different ranks and number of singularities, all isomorphic to each other. In each representation, the vertices of the core diagram are in one-to-one correspondence with the exponential factors of the connections. There is one generic representation, of maximal rank, with only one (irregular) singularity, located at infinity, and every $i=1,\dots,k$, defines a nongeneric representation, in which the vertices of the $i$-th subset of the core diagram correspond to regular singularities at finite distance, while all other core vertices correspond to exponential factors at infinity. In other words, one can read from the diagram $k$+1 representations of the same de Rham moduli space (see the example on the figure below). More concretely, the generic representation corresponds to connections of the form
$d-(Az+B)dz,$
with $A,B$ constant matrices, and $A$ semisimple, with $k$ distinct eigenvalues in one-to-one correspondence with the subsets of the partition of the core diagram. Furthermore, the equations governing the isomonodromic deformations of each representation are equivalent, so we can say that the different representations correspond to isomorphisms of isomonodromy systems.
Figure 1. Example of different readings of a (core) diagram in the simply-laced case. On the left, a complete 3-partite graph, with the different colours of the vertices corresponding to the partition. The three other figures on the right correspond to the nongeneric readings, each reading consisting in singling out one of the susbsets of the partition.
1.3. Basic operations on connections in genus zero and weak representations
From a more abstract point of view, Harnad duality is underlied by the Fourier-Laplace transform [28, 5, 7, 36]. It is an operation on (almost all) irreducible algebraic connections on Zariski open subsets of the affine line induced by the automorphism of the Weyl algebra of differential operators $A_{1}:=\mathbb{C}[z,\partial_{z}]$ defined by $z\mapsto-\partial_{z}$, $\partial_{z}\mapsto z$. Applied to such a connection, it gives another connection with in general different rank, number of singularities and pole orders.
@@ 1.3. Basic operations on connections in genus zero and weak representations
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
in $\mathrm{SL}_{2}(\mathbb{C})$ defines an automorphism of the Weyl algebra by $z\mapsto az+b$, $\partial_{z}\mapsto cz+d$. This induces an action of the group $\mathrm{SL}_{2}(\mathbb{C})$ on the set of isomorphism classes of irreducible algebraic connections on Zariski open subsets of the affine line $\mathbb{C}$ (excluding rank 1 connections with only a pole at infinity of order less than 2). The property of being of simply-laced type is preserved by this action, and the $k$+1 readings of the supernova quiver in the simply-laced case correspond to the different singularity patterns that appear in a given orbit.
More generally, there is a larger class of basic operations on irreducible irregular connections on $\mathbb{P}^{1}$, consisting of:
@@
    Möbius transformations of $\mathbb{P}^{1}$;
@@
    Twists by rank one connections;
@@
We will often refer to the $\operatorname{SL}_{2}(\mathbb{C})$ transformations as symplectic transformations to avoid possible confusions with Möbius transformations. These operations play an essential role in the extension due to Deligne-Arinkin [17, 3] of the Katz algorithm for rigid local systems to the case of irregular connections, which provides a way to reduce any irreducible rigid connection on $\mathbb{P}^{1}$ to the trivial rank one connection by repeated application of such operations.
Möbius transformations and twists induce in a straightforward way isomorphisms of nonabelian Hodge spaces, in particular they do not change their Betti side description as wild character varieties. On the other hand, the Fourier-Laplace transform is expected to induce nontrivial isomorphisms of wild character varieties. This has been proven in some cases [45] but more work needs to be done for the general case, due to the fact that determining the Stokes data of the Fourier transform of a genus zero connection is quite involved [37, 29, 43, 23, 31, 38, 39, 19]. For this reason, since in the meantime we still want to view genus zero wild Riemann surfaces with boundary data related by basic operations as representations of the same nonabelian Hodge space, relying on the fact that basic operations on connections induce well-defined counterparts at the level of the corresponding wild Riemann surfaces with boundary data, we make the following definition:
@@
In the rest of the article, we will allow ourselves to simply use the term representation to refer to weak representations. Given a (weak) representation $\mathbf{\Sigma}$ of a nonabelian Hodge space $\mathcal{M}$, we say that the representations obtained from $\Sigma$ by the action of a symplectic transformation defined by an element of $\operatorname{SL}_{2}(\mathbb{C})$, are its nearby representations. Finally, we will call basic representations of a nonabelian Hodge space the nearby representations of a representation of minimal rank.
@@
1.4. Main results: from enriched trees to classes of nearby representations
@@
The main purpose of this article is to develop a combinatorial way to find many weak representations of any nonabelian Hodge space in genus zero. Namely, we generalise the picture of obtaining different representations of a nonabelian Hodge space via different readings of a diagram from the simply-laced case to the case of genus zero nonabelian Hodge spaces with arbitrary singularity data. More specifically, we define some combinatorial data generalising the supernova quivers of the simply-laced case, from which several (expected to be) isomorphic nonabelian Hodge spaces can be explicitly obtained, corresponding to the different types of singularities appearing in an arbitrary $\operatorname{SL}_{2}(\mathbb{C})$ orbit, i.e. to different types of nearby representations of a given wild nonabelian Hodge space.
@@
We build on our previous work [18], in which a construction of a diagram associated to an irregular connection on a Zariski open subset of $\mathbb{P}^{1}$ was formulated, for the general case of arbitrary irregular types. Furthermore it was shown that the dimension of the wild character variety is given by a formula involving the Cartan matrix of the diagram, in agreement with the philosophy of ‘Global Lie theory’ advocated in [9] suggesting to view nonabelian Hodge spaces as some global analogues of Lie groups. The main property of the diagram is that it is invariant under the symplectic $\operatorname{SL}_{2}(\mathbb{C})$ action, so that connections with different types of singularities have the same diagram. It turns out that, unlike for the simply-laced case, the diagram is not quite enough to reconstruct the corresponding nonabelian Hodge spaces, so some extra data will be needed. In particular, the core diagram is not a complete $k$-partite graph any more in the general case, and these extra data will include a canonical partition of the core diagram.
@@
To obtain the appropriate extra data, as well as to give a precise meaning to the idea of ‘different types’ of singularity data, as was already sketched in [18, §5.3], we rely on our recent study of admissible deformations of wild Riemann surfaces [21, 20, 11, 22]. Being admissible deformations of each other defines an equivalence relation on the set of all possible wild Riemann surfaces of such connections, such that up to isomorphism the wild character variety $\mathcal{M}_{B}(\Sigma,\mathbf{a},\bm{\Theta},\bm{\mathcal{C}})$ is invariant under admissible deformations. This motivates the following definition:
@@
In [11], we showed that a class of admissible deformations of wild Riemann surfaces is completely characterised by a pair $(g,\mathbf{F})$, where $g$ is the genus of the underlying Riemann surface $\Sigma$, $\mathbf{F}$ is a fission forest, i.e. a collection of fission trees (see [11, Def. 3.18]). We can thus write $\mathcal{M}=\mathcal{M}(g,\mathbf{F},\bm{\mathcal{C}})$ up to isomorphism, where $\bm{\mathcal{C}}$ is the datum of a conjugacy class in $\operatorname{GL}_{n}(\mathbb{C})$ for each leaf of some fission tree of $\mathbf{F}$ (with $n\geq 1$ an integer, the multiplicity of the leaf). Here we only consider the case $g=0$, so we will drop the genus from the notation and write $\mathcal{M}=\mathcal{M}(\mathbf{F},\bm{\mathcal{C}})$.
The data that will allow us to reconstruct the classes of nearby representations consist of an enriched tree, defined as follows:
@@
    $\mathbf{T}$ is a short fission tree of , i.e. a fission tree such that all its vertices have height $\leq 2$.
@@
    If $N$ denotes the set of leaves of $\mathbf{T}$, $\bm{\mathcal{C}}$ is the datum of a conjugacy class $\mathcal{C}_{i}\in\operatorname{GL}_{n_{i}}(\mathbb{C})$, for some integer $n_{i}\geq 1$, for every $i\in N$.
@@
We show that enriched trees provide an invariant of irregular connections on $\mathbb{P}^{1}$ under symplectic transformations:
Theorem 1.4 (Prop. 3.11, Def. 3.12, Cor. 3.13).
@@
Any algebraic connection $(E,\nabla)$ on a Zariski open subset of $\mathbb{P}^{1}$ canonically determines an enriched tree $\mathscr{T}(E,\nabla)$, which depends only on its wild Riemann surface with boundary data. If $(E,\nabla)$ is irreducible, and is not a rank one connection with a pole of order less than 2 at infinity, and $A\in\operatorname{SL}_{2}(\mathbb{C})$, $A\cdot(E,\nabla)$ is well-defined and we have
$\mathscr{T}(A\cdot(E,\nabla))=\mathscr{T}(E,\nabla).$
@@
Our main result is that this invariant is powerful enough to allow us to reconstruct all classes of nearby representations:
Theorem 1.5 (Cor. 3.10).
@@
Let $\mathscr{T}=(\mathbf{T},\bm{\mathcal{C}})$ be an enriched tree, and $k$ be the number of principal subtrees of $\mathbf{T}$, i.e. the number of vertices of height 2. Assuming that the nonabelian Hodge space ($\mathcal{M}(\{\mathbf{T}\},\bm{\mathcal{C}})$ is nonempty, $\mathscr{T}$ canonically determines in an explicit way $k$ classes of representations $(\mathbf{F}_{i},\bm{\mathcal{C}_{i}})$, for $i=1,\dots,k$, of ($\mathcal{M}(\{\mathbf{T}\},\bm{\mathcal{C}})$, such that, if $(\bm{\Theta},\bm{\mathcal{C}})$ is any wild Riemann surface with boundary data on $\mathbb{P}^{1}$ with enriched tree $(\mathbf{T},\bm{\mathcal{C}})$, then ($(\{\mathbf{T}\},\bm{\mathcal{C}})$ and $(\mathbf{F}_{i},\bm{\mathcal{C}_{i}})$, for $i=1,\dots,k$ are all the classes of its nearby representations of ($\mathcal{M}(\{\mathbf{T}\},\bm{\mathcal{C}})$.

### v3

@@
[…]
Basic Representations of Genus Zero Nonabelian Hodge Spaces
[…]
Basic Representations of Genus Zero Nonabelian
Hodge Spaces
[…]
Jean DOUÇOT
[…]
J. Douçot
[…]
“Simion Stoilow” Institute of Mathematics of the Romanian Academy,
[…]
Received May 28, 2025, in final form February 13, 2026; Published online March 20, 2026
[…]
@@
[…]
irregular connections; Lax representations; Painlevé equations; Fourier transform; nonabelian Hodge diagrams
[…]
34M40; 44A10; 32C38; 32G34
1 Introduction
1.1 General motivation: classification of (wild) nonabelian Hodge spaces
@@
The term ‘nonabelian Hodge space’, introduced in [11, Definition 7], refers to a class of manifolds that can be seen as moduli spaces for different types of objects. On one hand, they are moduli spaces of meromorphic connections with possibly irregular singularities on curves (the de Rham side). Via the Riemann–Hilbert–Birkhoff correspondence, they are also moduli spaces of generalised monodromy data, known as wild character varieties [9, 15] (the Betti side), and, via the wild nonabelian Hodge correspondence [5], as moduli spaces of irregular Higgs bundles (the Dolbeault side). Furthermore, they carry rich (symplectic, hyperkähler) geometric structures, and give rise to (isospectral and isomonodromic) integrable systems, including many classical integrable systems of interest in mathematical physics such as all Painlevé equations (see, e.g., [11] for an overview).
@@
A nonabelian Hodge space $\mathcal{M}$ depends on the choice of a tuple $(\Sigma,\mathbf{a},\bm{\Theta},\bm{\mathcal{C}})$, that we call a wild Riemann surface with boundary data. Here the triple $(\Sigma,\mathbf{a},\bm{\Theta})$ is a wild Riemann surface [9, 12], meaning that $\Sigma$ is a compact Riemann surface, $\mathbf{a}=\{a_{1},\dots,a_{m}\}\subset\Sigma$ is a finite set of points on $\Sigma$ corresponding to the singularities of the connections giving rise to points of $\mathcal{M}_{\rm dR}$, the global irregular class $\bm{\Theta}$ encodes the type of singularities that these connections have at each singular point, and $\bm{\mathcal{C}}$ is a collection of conjugacy classes, encoding the formal monodromies of the connections.
@@
Remarkably, it turns out that there exist some isomorphisms between moduli spaces defined by different wild Riemann surfaces with boundary data, corresponding to connections with different ranks, number of singularities and pole orders. It is expected that when this happens the moduli spaces are really isomorphic as nonabelian Hodge spaces, i.e., all the structures of the nonabelian Hodge package match in the natural way. In other words, we would like to view those different wild Riemann surfaces with boundary data as different realizations, or representations, of one single abstract nonabelian Hodge space $\mathcal{M}$, which raises the question of their classification. Notice that here we will only be considering representations coming from the case of vector bundles, however wild nonabelian Hodge spaces arising from vector bundles may also be isomorphic to nonabelian Hodge spaces arising from principal $G$-bundles, for $G$ a complex reductive group different from $\operatorname{GL}_{n}(\mathbb{C})$, see, for example, [14].
@@
Varying the wild Riemann surface, one can consider moduli spaces of irregular connections in families, giving rise to isomonodromic deformations [9]. If the connections are written down explicitly in coordinates, this gives systems of nonlinear PDEs which include, among the simplest nontrivial examples, all Painlevé equations [35] (see also [4, 28, 42] for works investigating more systematically the question of obtaining explicit Hamiltonians for isomonodromic deformations). In this picture, the wild Riemann surface $(\Sigma,\mathbf{a},\bm{\Theta})$ giving rise to a given isomonodromy system corresponds to what is called a Lax representation, Lax pair, or linearisation for the system in the literature on integrable systems, albeit in a more abstract sense than is often the case (typically the term Lax pair is used to refer to an explicit parametrisation in coordinates of the coefficients of the connections). The existence of nontrivial isomorphisms between different nonabelian Hodge spaces is related to the existence of different Lax representations for many Painlevé-type equations, and the question of the classification of nonabelian Hodge spaces extends to the problem of finding all possible Lax representations for a given isomonodromy system.
1.2 Diagrams and isomorphisms between moduli spaces of connections
@@
Interestingly, it turns out that in many genus zero cases, i.e., when the underlying Riemann surface is $\mathbb{P}^{1}(\mathbb{C})$, there is a clean combinatorial way to pass between different representations.
@@
Possibly the simplest instance where such “dualities” between different moduli spaces of connections on $\mathbb{P}^{1}$ occur is Harnad duality [1, 2, 30, 52], which relates dual connections on trivial vector bundles over the affine line of the form ${\rm d}-\bigl(A+P(Y-z)^{-1}Q\bigr){\rm d}z$ and ${\rm d}+\bigl(Y+Q(A-z)^{-1}P\bigr){\rm d}z$, where $P\colon W\to V$ and $Q\colon V\to W$ are linear maps, where $V$, $W$ are finite dimensional complex vector spaces, and $A\in\operatorname{End}(V)$, $Y\in\operatorname{End}(W)$. The connections on both sides have an irregular singularity at infinity, with a pole of order two, as well as regular singularities at finite distance, whose positions are given by the eigenvalues of $A$ and $Y$. Harnad showed that the isomonodromic deformations of these two connections, governed by the JMMS equations [34], are equivalent. The duality induces an isomorphism between the de Rham moduli spaces corresponding to both sides.
@@
Harnad duality can be conveniently described in a graphical way as follows. The linear maps $A$, $Y$, $P$, $Q$ can be viewed as representations of the quiver drawn below, and the duality corresponds to exchanging the roles of the two vertices:
VVWWAAYYPPQQ
@@
This multiduality can also be conveniently encoded in a graphical way, in terms of a supernova quiver: it consists of a core, which is a complete $k$-partite graph for some integer $k$ (i.e., the core is a disjoint union of $k$ subsets of vertices, and each vertex in one subset is linked by one edge to every vertex belonging to a different subset) to which are then glued legs, each leg being a linear (i.e., type A) quiver.
The main statement is then as follows [8]: such a supernova quiver, together with some extra data, canonically determines $k+1$ moduli spaces of connections of simply-laced type, with in general different ranks and number of singularities, all isomorphic to each other. In each representation, the vertices of the core diagram are in one-to-one correspondence with the exponential factors of the connections. There is one generic representation, of maximal rank, with only one (irregular) singularity, located at infinity, and every $i\in\{1,\dots,k\}$ defines a nongeneric representation, in which the vertices of the $i$-th subset of the core diagram correspond to regular singularities at finite distance, while all other core vertices correspond to exponential factors at infinity. In other words, one can read from the diagram $k+1$ representations of the same de Rham moduli space (see the example on the figure below). More concretely, the generic representation corresponds to connections of the form ${\rm d}-(Az+B){\rm d}z$, with $A$, $B$ constant matrices, and $A$ semisimple, with $k$ distinct eigenvalues in one-to-one correspondence with the subsets of the partition of the core diagram. Furthermore, the equations governing the isomonodromic deformations of each representation are equivalent, so we can say that the different representations correspond to isomorphisms of isomonodromy systems.
Figure 1: Example of different readings of a (core) diagram in the simply-laced case, cf. [6, p. 11]. On the left, a complete 3-partite graph, with the different colours of the vertices corresponding to the 3 subsets of the partition. The generic reading corresponds to rank 5 connections with one pole of order 3. The three other figures on the right correspond to the nongeneric readings, each reading being obtained by singling out one of the subsets of the partition and interpreting as coming from regular singularities at finite distance. The nongeneric readings correspond to connections with ranks 3, 3, 4, and pole orders 3+1+1{3+1+1}, $3+1+1$, $3+1$ respectively from left to right.
1.3 Basic operations on connections on $\bm{\mathbb{P}^{1}}$ and weak representations
[…]
@@ 1.3 Basic operations on connections on $\bm{\mathbb{P}^{1}}$ and weak representations
$A=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)$
in $\mathrm{SL}_{2}(\mathbb{C})$ defines an automorphism of the Weyl algebra by $z\mapsto az+b\partial_{z}$, $\partial_{z}\mapsto cz+d\partial_{z}$ (this simply follows from the fact that this transformation preserves the defining commutation relation $\partial_{z},z]=1$ of $A_{1}$). This induces an action of the group $\mathrm{SL}_{2}(\mathbb{C})$ on the set of isomorphism classes of irreducible algebraic connections on Zariski open subsets of the affine line $\mathbb{C}$ (excluding rank 1 connections with only a pole at infinity of order less than 2). The property of being of simply-laced type is preserved by this action, and the $k+1$ readings of the supernova quiver in the simply-laced case correspond to the different singularity patterns that appear in a given orbit.
More generally, there is a larger class of basic operations on irreducible irregular connections on $\mathbb{P}^{1}$, consisting of
@@
    Möbius transformations of $\mathbb{P}^{1}$,
@@
    Twists by rank one connections,
@@
We will often refer to the $\operatorname{SL}_{2}(\mathbb{C})$ transformations as symplectic transformations to avoid possible confusions with Möbius transformations. These operations play an essential role in the extension due to Deligne–Arinkin [3, 19] of the Katz algorithm for rigid local systems [38] to the case of irregular connections, which provides a way to reduce any irreducible rigid connection on $\mathbb{P}^{1}$ to the trivial rank one connection by repeated application of such operations.
Möbius transformations and twists induce in a straightforward way isomorphisms of nonabelian Hodge spaces, in particular they do not change their Betti side description as wild character varieties. On the other hand, the Fourier transform is expected to induce nontrivial isomorphisms of wild character varieties.
It is known in the case of regular singularities and in the simply-laced case that the wild character varieties on both sides of the Fourier transform are isomorphic [10]. In the case of regular singularities it has also been shown that the Fourier transform induces a hyperkähler isometry [50] between the nonabelian Hodge moduli spaces. Some work remains to be done to deal with the general case, due to the fact that determining the Stokes data of the Fourier transform of an irregular connection on $\mathbb{P}^{1}$ is quite involved [18, 21, 31, 33, 41, 43, 44, 48]. For this reason, since in the meantime we still want to view genus zero wild Riemann surfaces with boundary data related by basic operations as representations of the same nonabelian Hodge space, relying on the fact that basic operations on connections induce well-defined counterparts at the level of the corresponding wild Riemann surfaces with boundary data, we make the following definition.
@@
In the rest of the article, we will allow ourselves to simply use the term representation to refer to weak representations. Given a (weak) representation $\mathbf{\Sigma}$ of a nonabelian Hodge space $\mathcal{M}$, we say that the representations obtained from $\bm{\Sigma}$ by the action of a symplectic transformation defined by an element of $\operatorname{SL}_{2}(\mathbb{C})$, are its nearby representations. Finally, we will call basic representations of a nonabelian Hodge space the nearby representations of a representation of minimal rank.
@@
1.4 Main results: from enriched trees to classes of nearby representations
@@
The main purpose of this article is to develop a combinatorial way to find many weak representations of any nonabelian Hodge space in genus zero. Namely, we generalise the picture of obtaining different representations of a nonabelian Hodge space via different readings of a diagram from the simply-laced case to the case of genus zero nonabelian Hodge spaces with arbitrary singularity data. More specifically, we define some combinatorial data generalising the supernova quivers of the simply-laced case, from which several (expected to be) isomorphic nonabelian Hodge spaces can be explicitly obtained, corresponding to the different types of singularities appearing in an arbitrary $\operatorname{SL}_{2}(\mathbb{C})$ orbit, i.e., to different types of nearby representations of a given wild nonabelian Hodge space.
@@
We build on our previous work [20], in which a construction of a diagram associated to an irregular connection on a Zariski open subset of $\mathbb{P}^{1}$ was formulated, for the general case of arbitrary irregular types. Furthermore it was shown that the dimension of the wild character variety is given by a formula involving the Cartan matrix of the diagram, in agreement with the philosophy of ‘global Lie theory’ advocated in [10] suggesting to view nonabelian Hodge spaces as some global analogues of Lie groups. The main property of the diagram is that it is invariant under the symplectic $\operatorname{SL}_{2}(\mathbb{C})$ action, so that connections with different types of singularities have the same diagram.
However, it turns out that, unlike for the simply-laced case, the diagram is not quite enough to reconstruct the corresponding nonabelian Hodge spaces, so some extra data will be needed. In particular, the core diagram is not a complete $k$-partite graph any more in the general case, and these extra data will include a canonical partition of the core diagram.
@@
To obtain the appropriate extra data, as well as to give a precise meaning to the idea of ‘different types’ of singularity data, as was already sketched in [20, Section 5.3], we rely on our recent study of admissible deformations of wild Riemann surfaces [13, 22, 23, 24]. Being admissible deformations of each other defines an equivalence relation on the set of all possible wild Riemann surfaces with boundary data of such connections, such that up to isomorphism the wild character variety $\mathcal{M}_{B}(\Sigma,\mathbf{a},\bm{\Theta},\bm{\mathcal{C}})$ is invariant under admissible deformations. This motivates the following definition.
@@
In [13], we showed that a class of admissible deformations of wild Riemann surfaces is completely characterised by a pair $(g,\mathbf{F})$, where $g$ is the genus of the underlying Riemann surface $\Sigma$$\mathbf{F}$ is a fission forest, i.e., a collection of fission trees (see [13, Definition 3.18]). We can thus write $\mathcal{M}=\mathcal{M}(g,\mathbf{F},\bm{\mathcal{C}})$} up to isomorphism, where $\bm{\mathcal{C}}$ is the datum of a conjugacy class in $\operatorname{GL}_{n}(\mathbb{C})$ for each leaf of some fission tree of $\mathbf{F}$, with $n\geq 1$ an integer, the multiplicity of the leaf. Here we only consider the case $g=0$, so we will drop the genus from the notation and write $\mathcal{M}=\mathcal{M}(\mathbf{F},\bm{\mathcal{C}})$.
The data that will allow us to reconstruct the classes of nearby representations consist of an enriched tree. To explain what it is, recall from [13, Section 3.4] that a fission tree is a tuple $\mathbf{T}=(\mathcal{T},\mathbb{V},\mathbb{A},\mathbb{L},h,n)$} where $\mathcal{T}$ is a metrised tree, $\mathbb{V}\subset\mathcal{T}$ is a subset whose points are the vertices of the tree, $\mathbb{A}$ and $\mathbb{L}$ are two distinguished subsets of $\mathbb{V}$ such that $\mathbb{L}\subset\mathbb{A}$, $h\colon\mathcal{T}\to\mathbb{R}_{\geq 0}$ is a function that we call the heightheight, such that for any vertex $v\in\mathbb{V}$, $h(v)\in\mathbb{Q}$, and $h^{-1}(0)=:N$ is the set of leaves of the tree, and $n\colon N\to\mathbb{Z}_{\geq 1}$ consists of the datum of an integer multiplicity $n_{i}\geq 1$} for each leaf $i$ (see Section 2.2 for more details on fission trees).
@@
    $\mathbf{T}=(\mathcal{T},\mathbb{V},\mathbb{A},\mathbb{L},h,n)$ is a short fission tree, i.e., a truncated fission tree in the sense of [13, Section 3.5] such that all its vertices different from the root have height $\leq 2$.
@@
    If $N$ denotes the set of leaves of $\mathbf{T}$, $\bm{\mathcal{C}}$ is the datum of a conjugacy class $\mathcal{C}_{i}\in\operatorname{GL}_{n_{i}}(\mathbb{C})$ for each $i\in N$, where $n_{i}\geq 1$ is the multiplicity of the leaf $i$.
@@
We show that enriched trees provide an invariant of irregular connections on $\mathbb{P}^{1}$ under symplectic transformations.
Theorem 1.4 (Proposition 3.12, Definition 3.13, Corollary 3.14).
@@
Any algebraic connection $(E,\nabla)$ on a Zariski open subset of $\mathbb{P}^{1}$ canonically determines an enriched tree $\mathscr{T}(E,\nabla)$, which depends only on its wild Riemann surface with boundary data. If $(E,\nabla)$ is irreducible, and is not a rank one connection with a pole of order less than $2$ at infinity, then for $A\in\operatorname{SL}_{2}(\mathbb{C})$, $A\cdot(E,\nabla)$ is well-defined and we have $\mathscr{T}(A\cdot(E,\nabla))=\mathscr{T}(E,\nabla)$.
@@
Our main result is that this invariant is powerful enough to allow us to reconstruct all classes of nearby representations.
Theorem 1.5 (Corollary 3.11).
@@
Let $\mathscr{T}=(\mathbf{T},\bm{\mathcal{C}})$ be an enriched tree, and $k$ be the number of principal subtrees of $\mathbf{T}$, i.e., the number of vertices of height $2$. Assuming that the nonabelian Hodge space $\mathcal{M}(\{\mathbf{T}\},\bm{\mathcal{C}})$ is nonempty, $\mathscr{T}$ canonically determines in an explicit way $k$ classes of representations $(\mathbf{F}_{i},\bm{\mathcal{C}_{i}})$, for $i\in\{1,\dots,k\}$, of $\mathcal{M}(\{\mathbf{T}\},\bm{\mathcal{C}})$, such that, if $(\bm{\Theta},\bm{\mathcal{C}})$ is any wild Riemann surface with boundary data on $\mathbb{P}^{1}$ with enriched tree $(\mathbf{T},\bm{\mathcal{C}})$, then $(\{\mathbf{T}\},\bm{\mathcal{C}})$ and $(\mathbf{F}_{i},\bm{\mathcal{C}_{i}})$, for $i\in\{1,\dots,k\}$ are all the classes of its nearby representations of $\mathcal{M}(\{\mathbf{T}\},\bm{\mathcal{C}})$.